OpenStudy (anonymous):

Find dy/dx y=(116-21x^2)/120 Hi! I dont understand what to do here, please help!

OpenStudy (anonymous):

\[y=\frac{ 116-21x^2 }{ 120 }\] here is the equation again

OpenStudy (seascorpion1):

\[y=\frac{116}{120}-\frac{21}{120}x^2\] therefore: \[y'=-\frac{21*2}{120}x^{2-1}\]\[=-\frac{21}{60}x\]

OpenStudy (dayakar):

\[y= \frac{ 116 }{ 120 }-\frac{ 21x^{2} }{ 120 }\]

OpenStudy (dayakar):

now find dy/dx

OpenStudy (anonymous):

the \[\frac{ 116 }{ 120 } \] goes to zero when applying the power rule (I think thats what its called) right? and then the next part goes to \[\frac{ -21 }{ 60 }x\] but the answer in my text book says its -7x/20, so you divide both sides by three to get that

OpenStudy (dayakar):

-21x/60 (cancellation with 3 ) =-7x/20

OpenStudy (dayakar):

u r right

OpenStudy (anonymous):

Ok great! Thanks! :D

OpenStudy (dayakar):

u r welcome:)